OpenStudy (anonymous):

A rain gutter is made from sheets of aluminum that are 20 inches wide by turning up the edges to form right angles. Determine the depth of the gutter that will maximize its cross sectional area to allow the greatest amount of water to flow

OpenStudy (anonymous):

Perimeter of 3 sides: y + 2x = 20 in -> y = 20 - 2x Area : A = x ( 20 - 2x) = - 2x² + 20x Max Area: A' = - 4x + 20 = 0 --> x = 5 in